# Cancellation properties for exotic $4$-dimensional positive scalar curvature metrics

### v2

@@ Abstract.
More generally, we give easily checkable conditions on diffeomorphisms $f:M\to M$ of $1$-connected $4$-manifolds which guarantee that for each closed $N$ and each $g\in\mathcal{R}^{+}(M)$, the metrics $\nu_{N}(g)$ and $\nu_{N}(f^{*}g)$ lie in the same path component of $\mathcal{R}^{+}(M\times N)$. The conditions are (A) the cobordism class of the mapping torus $[T(f)]\in\Omega_{5}^{\mathrm{SO}}$ vanishes and (B) the induced map $f^{*}:H^{2}(M;\mathbb{R})\to H^{2}(M;\mathbb{R})$ belongs to the unit component of $\mathrm{Aut}(I_{M})$, the orthogonal group of the intersection form of $M$. Depending on whether $\dim(N)=1$ or $\dim(N)\geq 2$ and on whether or not $M$ is spin, various combinations of (A) and (B) are needed.
@@ 1. Introduction
Let $M_{0}$ and $M_{1}$ be closed, $1$-connected smooth $4$-manifolds with isomorphic intersection form. Then
@@
    $M_{0}$ and $M_{1}$ are $h$-cobordant (Wall [42, Theorem 2]),
@@
    we have handle stabilization: $M_{0}\sharp k(S^{2}\times S^{2})$ and $M_{1}\sharp k(S^{2}\times S^{2})$ are diffeomorphic for sufficiently large $k$ (Wall [42, Theorem 3]),
@@
It is well-known that in the high-dimensional regime ($\dim(M)\geq 5$), the theory of positive scalar curvature is governed by cobordism theory; this is due to the Gromov–Lawson surgery theorem [21] [6]; see [11, Theorems 1.2 and 1.5] for the most general formulation of the cobordism theorem. In dimension $4$, we can at least infer that this is true “stably”, in the following sense.
@@
Let $B\stackrel{{\scriptstyle\theta}}{{\to}}BO$ be a fibration and let $M_{0}$ be a closed $4$-manifold, together with a lift $\ell:M_{0}\to BO$ of the stabilized tangent bundle of $M_{0}$, and suppose that $\ell$ is $2$-connected.
Assume furthermore that the cobordism class $[M_{0}]\in\Omega_{4}^{\theta}$ contains a representative $[M_{1}]$ which has a psc metric. Then
@@
This result is certainly known and a rather easy consequence of the Gromov–Lawson surgery theorem. We give the proof in §2. If $M_{0}$ had dimension $\geq 5$, the conclusion that $M_{0}$ has a psc metric is valid without stabilization, by [11, Theorem 1.5]. The latter conclusion fails in dimension $4$:
@@
Now a $4$-dimensional Kähler manifold $M$ with $b_{2}^{+}(M)\geq 2$ does not have a psc metric, by a result of Taubes [41], combined with e.g. [33, Corollary 5.18].
Examples of such Kähler manifolds which are $1$-connected and non-spin are easy to find: a degree $d$ algebraic hypersurface $X_{d}\subset\mathbb{C}\mathbb{P}^{3}$ is $1$-connected by the Lefschetz hyperplane theorem, and its characteristic numbers can be calculated by a classical method [31]. If $d\geq 5$ is odd, $X_{d}$ is nonspin and has $b_{2}^{+}(X_{d})\geq 2$.
Let us now turn, finally, to the main aim of this note, that is $\pi_{0}(\mathcal{R}^{+}(M))$. There are examples of $4$-manifolds with “exotic” $\pi_{0}(\mathcal{R}^{+}(\_))$. For instance, the following is proven in [37, §5].
@@
There exist a $1$-connected $4$-manifold [^1] For example $M=4(\mathbb{C}\mathbb{P}^{2})\sharp 21(\overline{\mathbb{C}\mathbb{P}^{2}})$. $M$, a psc metric $g_{0}\in\mathcal{R}^{+}(M)$ and a family $\{f_{n}|n\in\mathbb{N}_{0}\}$ of orientation-preserving diffeomorphisms $f_{n}:M\to M$ with $f_{0}=\mathrm{id}$, with the following properties.
@@
    Each $f_{n}$ is pseudoisotopic to the identity, i.e. there are diffeomorphisms $F_{n}:M\times[0,1]\to M\times[0,1]$ such that $F_{n}(x,0)=(x,0)$ and $F_{n}(x,1)=(f_{n}(x),1)$ for each $x\in M$.
@@
    The metrics $g_{n}:=f_{n}^{*}g_{0}$ are pairwise nonisotopic, i.e. the elements $[g_{n}]\in\pi_{0}(\mathcal{R}^{+}(M))$ are pairwise distinct.
Not surprisingly, that $[g_{n}]\neq[g_{m}]$ for $n\geq m$ is detected by Seiberg–Witten invariants. Note that the psc metrics $g_{n}$ are all concordant, via $F_{n}^{*}(g_{0}\oplus dx^{2})$. To best of the knowledge of the author, these are the only known examples of concordant but not isotopic psc metrics (index theoretic methods cannot address this question at all, see [24] for a high-brown variant of this statement).
These examples satisfy “handle stabilization”: after taking connected sums with sufficiently many $S^{2}\times S^{2}$’s, all the metrics $g_{n}$ become isotopic. More precisely, we may suppose that there is an embedded disc $D^{4}\subset M$ fixed (pointwise) by each $f_{n}$, and we may also suppose that $g_{0}$ is a torpedo metric inside $D^{4}$. We can therefore take the connected sum of $f_{n}$ and the identity on $k(S^{2}\times S^{2})$, and likewise we can take the connected sum of $g_{0}$ with some suitable standard metric on $k(S^{2}\times S^{2})$. It follows from [36, Theorem 1.4] (and correction [17]) that after taking connected sum with $k(S^{2}\times S^{2})$ for suitable $k$, $f_{n}$ becomes isotopic to the identity, so that $g_{n}$ and $g_{0}$ become isotopic (handle stabilization).
Note that for $M^{2n}$ with $n\geq 3$, the connected sum map $\mathcal{R}^{+}(M)\to\mathcal{R}^{+}(M\sharp(S^{n}\times S^{n}))$ is a homotopy equivalence, by [11, Theorem 1.2]; by Theorem 1.2, this is not true for $n=2$.
We wish to prove that in the situation of Theorem 1.2, we also have product stabilization, but beforehand have to clarify what is meant by that.
@@
A very easy scaling argument shows that the inclusion $\mathcal{R}^{>\epsilon}(M)\to\mathcal{R}^{+}(M)$ is a homotopy equivalence, whereas $\mathcal{R}^{>-\epsilon}(N)$ is contractible by [14, Lemma 5.4]. Hence we may interprete $\nu$, up to homotopy, as a map
@@
The homotopical behavior of $\nu_{N}$ depends strongly on $N$. If $N$ does admit a metric of positive scalar curvature, the map $\nu_{N}$ factors through some $\mathcal{R}^{-\epsilon}(M)$ and is therefore homotopic to a constant.
@@
[…]
commutes; here $\alpha(N)\in KO_{n}(C^{*}(G))$ is the Rosenberg index. If e.g. $N=S^{1}$, the right vertical map is injective and if $d\geq 6$, the top horizontal map is very much nontrivial by the main result of [4], so $(\nu_{N})_{*}$ is nontrivial.
Let us turn to the $4$-dimensional world. In light of the construction of the metrics $g_{n}$, it seems plausible that they also exhibit product stabilization. Our main aim is to confirm that this is indeed the case:
@@
Let $M$ and $g_{n}:=f_{n}^{*}g_{0}$ be as in Theorem 1.2 and let $N$ be a closed connected smooth manifold of positive dimension. Then the metrics $\nu_{N}(g_{n})$ all lie in the same path component of $\mathcal{R}^{+}(M\times N)$.
As we said above, this is trivially true if $N$ has a psc metric. For $n\geq 3$, we could argue by pseudoisotopy theory (e.g. [23, p. 11] as follows. Since $\pi_{1}(M)=1$, the pseudoisotopy obtruction of $f_{n}$ vanishes. Therefore, so does that of $f_{n}\times\mathrm{id}_{N}$, and if $n\geq 3$, the quoted result implies that $f_{n}\times\mathrm{id}_{N}$ is isotopic to the identity. In light of the construction of $g_{n}$, this implies Theorem 1.6 for $\dim(N)\geq 3$ (the author was informed by Daniel Ruberman that a direct proof that $f_{n}\times\mathrm{id}_{S^{1}}$ is isotopic to the identity is possible, which also gives a proof of Theorem 1.6 for $\dim(N)=1$).
The purpose of this note is to give a proof of Theorem 1.6 that also works for $\dim(N)=1,2$ is cobordism-theoretic in nature rather than through pseudoisotopy, and does not depend on an understanding how the diffeomorphisms $f_{n}$ are constructed. Moreover, we only need very mild homological conditions on $f_{n}$ and not that $f_{n}$ is pseudoisotopic to the identity.
For the formulation of the result, recall that the mapping torus $T(f)$ of $f\in\mathrm{Diff}(M)$ is $T(f):=(M\times[0,1])/(x,1)\sim(f(x),0)$, which is a fibre bundle over $S^{1}$ with fibre $M$. If $M$ is oriented and $f$ is orientation–preserving, $T(f)$ has a natural orientation. Assigning to $f$ the cobordism class $[T(f)]\in\Omega_{d+1}^{\mathrm{SO}}$ ($d=\dim(M)$) gives a map
@@
which is a group homomorphism. We remark that $\Omega^{\mathrm{SO}}_{5}\cong\mathbb{Z}/2$, via the Stiefel–Whitney number $[M]\mapsto w_{2}w_{3}(M)$, and that $w_{2}w_{3}(T(f))$ can be computed in terms of the action of $f$ on $H^{*}(M;\mathbb{R})$ and $H^{*}(M;\mathbb{F}_{2})$:
@@
(see Lemma 4.7 below).
Moreover, let $I_{M}:H^{2}(M;\mathbb{R})\times H^{2}(M;\mathbb{R})\to\mathbb{R}$ be the intersection form of $M$, a symmetric nondegenerate bilinear form on a finite-dimensional $\mathbb{R}$-vector space. We let $\mathrm{Aut}(I_{M})\subset\mathrm{GL}(H^{2}(M;\mathbb{R}))$ be the orthogonal group of $I_{M}$. For $f\in\mathrm{Diff}^{+}(M)$, the induced map $H^{2}(f;\mathbb{R})$ belongs to $\mathrm{Aut}(I_{M})$. We view $\mathrm{Aut}(I_{M})$ as a Lie group; as such it is isomorphic to the indefinite orthogonal group $O(b_{2}^{+}(M),b_{2}^{-}(M))$. It is well-known (and will be recalled later) that $\pi_{0}(\mathrm{Aut}I_{M})\cong\mathbb{Z}/2$ or $\cong(\mathbb{Z}/2)^{2}$, depending on whether $I_{M}$ is definite or not.
@@
    $n\geq 2$, $M$ is not spin and $[T(f)]=0\in\Omega^{\mathrm{SO}}_{5}$;
@@
    $n=1$, $M$ is spin, and $[H^{2}(f;\mathbb{R})]=1\in\pi_{0}(\mathrm{Aut}(I_{M}))$;
@@
    $n=1$, $M$ is not spin, $[T(f)]=0$ and $[H^{2}(f;\mathbb{R})]=1\in\pi_{0}(\mathrm{Aut}(I_{M}))$.
The proof rests on rigidity theorems for the action of diffeomorphism groups on $\mathcal{R}^{+}(M)$ for high-dimensional $M$, in particular on the works by Frenck [16] (when $n\geq 2$) and Bantje [2] (when $n=1$). We give a review of these results in §3. What these results have in common is that they govern the action of $\mathrm{Diff}(M)$ on $\mathcal{R}^{+}(M)$ in terms of cobordism theory; we give more precise statements below. Using these results, the proof of Theorem B comes down to verify that the mapping torus $T(f)$ vanishes in a certain cobordism group. When $n\geq 2$, the relevant cobordism group is just $\Omega^{\mathrm{Spin}}_{5}=0$ or $\Omega_{5}^{\mathrm{SO}}$. When $n=1$, the relevant cobordism group is the first homotopy group $\pi_{1}(\mathrm{MT}\mathrm{SO}(4))$ or $\pi_{1}(\mathrm{MT}\mathrm{Spin}(4))$ of the Madsen–Tillmann spectra. In §4, we give a detailed calculation of these groups. The answer we find accounts for the conditions of (3) and (4) of Theorem B.
Let us now turn to the last of our main results, concerning a recent result by Auckly and Ruberman [1, Theorem 1.14] that reads as follows.
@@
For each $p$ and $j\leq p$ even, there are manifolds $M$ of the form $M=k(S^{2}\times S^{2})$ or $M=k(\mathbb{C}\mathbb{P}^{2}\sharp\overline{\mathbb{C}\mathbb{P}^{2}})$ for large $k$ and $g_{0}\in\mathcal{R}^{+}(M)$ such that
@@
    $\ker(\pi_{j}(\mathrm{Diff}^{+}(M))\to\pi_{j}(\mathrm{Homeo}^{+}(M)))$ contains a free abelian group of infinite rank,
@@
    The map $\pi_{j}(\mathrm{Diff}^{+}(M))\to\pi_{j}(\mathcal{R}^{+}(M))$ induced by the orbit map $f\mapsto f^{*}g_{0}$, is injective on the free abelian subgroup of (1).
Here is the last of our results.
@@
Let $M$ and $j$ as in Theorem 1.9. For any $N$ of positive dimension and any $j>0$, the map
@@
kills all the elements of Theorem 1.9 (2).

### v3

@@ Abstract.
More generally, we give easily checkable conditions on diffeomorphisms $f:M\to M$ of $1$-connected 44-manifolds which guarantee that for each closed $N$ and each $g\in\mathcal{R}^{+}(M)$, the metrics $\nu_{N}(g)$ and $\nu_{N}(f^{*}g)$ lie in the same path component of $\mathcal{R}^{+}(M\times N)$. The conditions are (A) the cobordism class of the mapping torus $T(f)]\in\Omega_{5}^{\mathrm{SO}}$ vanishes and (B) the induced map $f^{*}:H^{2}(M;\mathbb{R})\to H^{2}(M;\mathbb{R})$ belongs to the unit component of $\mathrm{Aut}(I_{M})$, the orthogonal group of the intersection form of $M$. Depending on whether $\dim(N)=1$ or $\dim(N)\geq 2$ and on whether or not $M$ is spin, various combinations of (A) and (B) are needed.
@@ 1. Introduction
Let $M_{0}$ and $M_{1}$ be closed, $1$-connected smooth $4$-manifolds with isomorphic intersection forms. Then
@@
    $M_{0}$ and $M_{1}$ are $h$-cobordant (Wall [47, Theorem 2]),
@@
    we have handle stabilization: $M_{0}\sharp k(S^{2}\times S^{2})$ and $M_{1}\sharp k(S^{2}\times S^{2})$ are diffeomorphic for sufficiently large $k$ (Wall [47, Theorem 3]),
@@
It is well-known that in the high-dimensional regime ($\dim(M)\geq 5$), the theory of positive scalar curvature is governed by cobordism theory; this is due to the Gromov–Lawson surgery theorem [24] [9]; see [14, Theorems 1.2 and 1.5] for the most general formulation. In dimension $4$, we can at least infer that this is true “stably”, in the following sense.
@@
Let $B\stackrel{{\scriptstyle\theta}}{{\to}}BO$ be a fibration and let $M_{0}$ be a closed $4$-manifold, together with a lift $\ell:M_{0}\to B$ of the stabilized tangent bundle of $M_{0}$, and suppose that $\ell$ is $2$-connected. Assume furthermore that the cobordism class $M_{0}]\in\Omega_{4}^{\theta}$ contains a representative $M_{1}]$ which has a psc metric. Then
@@
This result is certainly known and a rather easy consequence of the Gromov–Lawson surgery theorem. We give the proof in §2. If $M_{0}$ had dimension $\geq 5$, the conclusion that $M_{0}$ has a psc metric is valid without stabilization, by [14, Theorem 1.5]. The latter conclusion fails in dimension $4$:
@@
A $4$-dimensional Kähler manifold $M$ with $b_{2}^{+}(M)\geq 2$ does not have a psc metric, by a result of Taubes [46], combined with e.g. [36, Corollary 5.18].
Examples of such Kähler manifolds which are $1$-connected and non-spin are easy to find: a degree $d$ algebraic hypersurface $X_{d}\subset\mathbb{C}\mathbb{P}^{3}$ is $1$-connected by the Lefschetz hyperplane theorem, and its characteristic numbers can be calculated by a classical method [34]. The outcome is that for $d\geq 5$ is odd, $X_{d}$ is nonspin and has $b_{2}^{+}(X_{d})\geq 2$.
Let us now turn, finally, to the main aim of this note, that is $\pi_{0}(\mathcal{R}^{+}(M))$. There are examples of $4$-manifolds with “exotic” $\pi_{0}(\mathcal{R}^{+}(\_))$. For instance, the following is proven in [40, §5].
@@
There exists a $1$-connected 44-manifold [^1] For example $4(\mathbb{C}\mathbb{P}^{2})\sharp 21(\overline{\mathbb{C}\mathbb{P}^{2}}). MM$, a psc metric $g_{0}\in\mathcal{R}^{+}(M)$ and a family $\{f_{i}|i\in\mathbb{N}_{0}\}$ of orientation-preserving diffeomorphisms $f_{i}:M\to M$ with $f_{0}=\mathrm{id}$, with the following properties.
@@
    Each $f_{i}$ is pseudoisotopic to the identity, i.e. there are diffeomorphisms $F_{i}:M\times[0,1]\to M\times[0,1]$ such that $F_{i}(x,0)=(x,0)$ and $F_{i}(x,1)=(f_{i}(x),1)$ for each $x\in M$.
@@
    The metrics $g_{i}:=f_{i}^{*}g_{0}$ are pairwise nonisotopic, i.e. the elements $g_{i}]\in\pi_{0}(\mathcal{R}^{+}(M))$ are pairwise distinct.
Not surprisingly, that $g_{i}]\neq[g_{j}]$ for $i\neq j$ is detected by Seiberg–Witten invariants. Note that the psc metrics $g_{i}$ are all concordant, via $F_{i}^{*}(g_{0}\oplus dx^{2})$. To best of the knowledge of the author, these are the only known examples of concordant but not isotopic psc metrics (index theoretic methods cannot address this question at all, see [27] for a high-brow variant of this statement).
These examples satisfy “handle stabilization”: after taking the connected sum with sufficiently many copies of $S^{2}\times S^{2}$’s, all the metrics $g_{i}$ become isotopic. More precisely, we may suppose that there is an embedded disc $D^{4}\subset M$ fixed (pointwise) by each $f_{i}$, and we may also suppose that $g_{0}$ is a torpedo metric inside $D^{4}$. We can therefore take the connected sum of $f_{i}$ and the identity on $k(S^{2}\times S^{2})$, and likewise we can take the connected sum of $g_{0}$ with some suitable standard metric on $k(S^{2}\times S^{2})$. It follows from [39, Theorem 1.4] (as corrected by [20]) that after taking connected sum with $k(S^{2}\times S^{2})$ for suitable $k$, $f_{i}$ becomes isotopic to the identity, so that $g_{i}$ and $g_{0}$ become isotopic (handle stabilization).
Note that for $M^{2n}$ with $n\geq 3$, the connected sum map
$\mathcal{R}^{+}(M)\to\mathcal{R}^{+}(M\sharp(S^{n}\times S^{n}))$
is a homotopy equivalence, by [14, Theorem 1.2]; the above reasoning proves that this is not true when $n=2$.
We wish to prove that in the situation of Theorem 1.2, we also have product stabilization. Beforehand, we have have to clarify what is meant by that.
@@
A very easy scaling argument shows that the inclusion $\mathcal{R}^{>\epsilon}(M)\to\mathcal{R}^{+}(M)$ is a homotopy equivalence, whereas $\mathcal{R}^{>-\epsilon}(N)$ is contractible by [17, Lemma 5.4]. Hence we may interpret $\nu$, up to homotopy, as a map
@@
The homotopical behavior of $\nu_{N}$ depends strongly on $N$.
  1. (1)
    If $N$ does admit a metric of positive scalar curvature, the map $\nu_{N}$ factors through some $\mathcal{R}^{-\epsilon}(M)$ and is therefore homotopic to a constant.
  2. (2)
@@
    ${\pi_{k}(\mathcal{R}^{+}(M))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(\nu_{N})_{*}}$$\scriptstyle{\mathrm{inddiff}}$${KO_{k+d+1}(*)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\_\times\hat{\mathfrak{a}}_{G}(N)}$${\pi_{k}(\mathcal{R}^{+}(M\times N))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathrm{inddiff}^{G}}$${KO_{k+d+n+1}(C^{*}(G))}$
    commutes; here $\hat{\mathfrak{a}}(N)\in KO_{n}(C^{*}(G))$ is the Rosenberg index of $N$. If e.g. $N=S^{1}$, the right vertical map is injective and if $d\geq 6$, the top horizontal map is very much nontrivial by the main result of [7], so $(\nu_{N})_{*}$ is nontrivial.
Let us turn to the $4$-dimensional world. In light of the construction of the metrics $g_{i}$ in Theorem 1.2, it seems plausible that they also exhibit product stabilization. Our main aim is to confirm that this is indeed the case:
@@
Let $M$ and $g_{i}:=f_{i}^{*}g_{0}$ be as in Theorem 1.2 and let $N$ be a closed connected smooth manifold of dimension $n\geq 1$. Then the metrics $\nu_{N}(g_{i})$ all lie in the same path component of $\mathcal{R}^{+}(M\times N)$.
By Remark 1.5, this is trivially true if $N$ has a psc metric. For $n\geq 3$, we could argue by pseudoisotopy theory (e.g. [26, p. 11]) as follows. Since $\pi_{1}(M)=1$, the pseudoisotopy obtruction of $f_{i}$ vanishes. Therefore, so does that of $f_{i}\times\mathrm{id}_{N}$, and if $n\geq 3$, the quoted result implies that $f_{i}\times\mathrm{id}_{N}$ is isotopic to the identity. In light of the construction of $g_{i}$, this implies Theorem 1.6 for $\dim(N)\geq 3$ (the author was informed by Daniel Ruberman that a direct proof that $f_{i}\times\mathrm{id}_{S^{1}}$ is isotopic to the identity is possible, which also gives a proof of Theorem 1.6 for $\dim(N)=1$).
The purpose of this note is to give a proof of Theorem 1.6 that also works for $\dim(N)=1$ and $2$, is cobordism-theoretic in nature rather than through pseudoisotopy, and does not depend on an understanding how the diffeomorphisms $f_{i}$ are constructed. Moreover, we only need very mild homological conditions on $f_{i}$ and not that $f_{i}$ is pseudoisotopic to the identity.
For the formulation of the result, recall that the mapping torus $T(f)$ of $f\in\mathrm{Diff}(M)$ is $T(f):=(M\times[0,1])/(x,1)\sim(f(x),0)$, which is a fibre bundle over $S^{1}$ with fibre $M$. If $M$ is oriented and $f$ is orientation–preserving, $T(f)$ has a natural orientation. Assigning to $f$ the cobordism class $T(f)]\in\Omega_{d+1}^{\mathrm{SO}}$ ($d=\dim(M)$) gives a map
@@
which is a group homomorphism. We remark that $\Omega^{\mathrm{SO}}_{5}\cong\mathbb{Z}/2$, via the Stiefel–Whitney number $M]\mapsto w_{2}w_{3}(M)$. For mapping tori, $w_{2}w_{3}(T(f))$ can be computed in terms of the action of $f$ on $H^{*}(M;\mathbb{R})$ and $H^{*}(M;\mathbb{F}_{2})$ (see Lemma 4.7 below):
@@
(we use the notation $H^{k}(f;R)$ for the map induced by $f$ in cohomology with coefficients in $R$, and the notation $\mathrm{Eig}_{\lambda}(F)$ for the $\lambda$-eigenspace of a linear endomorphism $F$). Let $I_{M}:H^{2}(M;\mathbb{R})\times H^{2}(M;\mathbb{R})\to\mathbb{R}$ be the intersection form of $M$, a symmetric nondegenerate bilinear form on a finite-dimensional $\mathbb{R}$-vector space. We let $\mathrm{Aut}(I_{M})\subset\mathrm{GL}(H^{2}(M;\mathbb{R}))$ be the orthogonal group of $I_{M}$. For $f\in\mathrm{Diff}^{+}(M)$, the induced map $H^{2}(f;\mathbb{R})$ belongs to $\mathrm{Aut}(I_{M})$. We view $\mathrm{Aut}(I_{M})$ as a Lie group; as such it is isomorphic to the indefinite orthogonal group $O(b_{2}^{+}(M),b_{2}^{-}(M))$. It is well-known that $\pi_{0}(\mathrm{Aut}I_{M})\cong\mathbb{Z}/2$ or $\cong(\mathbb{Z}/2)^{2}$, depending on whether $I_{M}$ is definite or not.
@@
    $n\geq 2$, $M$ is not spin and $T(f)]=0\in\Omega^{\mathrm{SO}}_{5}$;
@@
    $n=1$, $M$ is spin, and $H^{2}(f;\mathbb{R})]=1\in\pi_{0}(\mathrm{Aut}(I_{M}))$;
@@
    $n=1$, $M$ is not spin, $[T(f)]=0$ and $H^{2}(f;\mathbb{R})]=1\in\pi_{0}(\mathrm{Aut}(I_{M}))$.
In Lemma 4.22, we give elementary examples of diffeomorphisms whose invariants in $\Omega^{\mathrm{SO}}_{5}$ and $\pi_{0}(\mathrm{Aut}(I_{M}))$ do not vanish.
The proof of the Theorem rests on rigidity theorems for the action of diffeomorphism groups on $\mathcal{R}^{+}(M)$ for high-dimensional $M$, in particular on the works by Frenck [19] (when $n\geq 2$) and Bantje [5] (when $n=1$). We give a review of these results in §3. What these results have in common is that they govern the action of $\mathrm{Diff}(M)$ on $\mathcal{R}^{+}(M)$ in terms of cobordism theory; we give more precise statements below. Using these results, the proof of Theorem B comes down to verifying that the mapping torus $T(f)$ vanishes in a certain cobordism group. When $n\geq 2$, the relevant cobordism group is just $\Omega_{5}^{\mathrm{SO}}$ or $\Omega^{\mathrm{Spin}}_{5}=0$. When $n=1$, the relevant cobordism group is the first homotopy group $\pi_{1}(\mathrm{MT}\mathrm{SO}(4))$ or $\pi_{1}(\mathrm{MT}\mathrm{Spin}(4))$ of the Madsen–Tillmann spectrum. In §4, we give a detailed calculation of these groups. The answer we find accounts for the conditions of (3) and (4) of Theorem B.
Let us now turn to the last of our main results, concerning a recent result by Auckly and Ruberman [4, Theorem 1.14] that reads as follows.
@@
For each $p\geq 0$ and $j\leq p$ with pjp-j even, there are manifolds $M$ of the form $M=k(S^{2}\times S^{2})$ or $M=k(\mathbb{C}\mathbb{P}^{2}\sharp\overline{\mathbb{C}\mathbb{P}^{2}})$ for large $k$ and $g_{0}\in\mathcal{R}^{+}(M)$ such that
@@
    $\ker(\pi_{j}(\mathrm{Diff}^{+}(M))\to\pi_{j}(\mathrm{Homeo}^{+}(M)))$ contains a free abelian group $A$ of infinite rank,
@@
    The map $o_{*}:\pi_{j}(\mathrm{Diff}^{+}(M))\to\pi_{j}(\mathcal{R}^{+}(M))$ induced by the orbit map $o:f\mapsto f^{*}g_{0}$, is injective on the free abelian subgroup $A$ of (1).
@@
Let $M$ and $j$ as in Theorem 1.9. For any NN of positive dimension and any $j>0$, the map
@@
vanishes on the subgroup $o_{*}(A)$ described in Theorem 1.9.